# Distinguishing 2-knots admitting circle actions by fundamental groups

### v3

@@ 1. Introduction
Circle actions on homotopy $4$-spheres are classified by Montgomery and Yang [24] and Fintushel [10]. Later, Pao proved in [26] that such homotopy $4$-spheres are the standard $4$-sphere $S^{4}$. A $2$-sphere embedded in $S^{4}$ that is invariant under a circle action is called a branched twist spin (cf. [18, §16.3]). Note that a circle embedded in $S^{3}$ that is invariant under a circle action is called a torus knot. A twist spun knot is an example of a branched twist spin. See Section 2.1 for more precise explanation. In this sense, branched twist spins are the simplest class of $2$-knots in $S^{4}$ if one wants to understand $2$-knots geometrically. The orbit space of a circle action on $S^{4}$ is the $3$-sphere $S^{3}$ except a trivial case and the image of the union of singular orbits by the orbit map constitutes an embedded circle, that is a $1$-knot, in $S^{3}$. Conversely, we may recover a branched twist spin from a $1$-knot in $S^{3}$ with a pair of coprime integers $m$ and $n$ with $n>0$ uniquely. For a given knot $K$ in $S^{3}$ and these integers $m$ and $n$, we denote the corresponding branched twist spin by $K^{m,n}$.
@@
The aim of this paper is to understand if we can obtain the same branched twist spin from different $1$-knots in $S^{3}$. This problem is studied by the first author in [11, 12, 14] using knot determinants, irreducible $SL(2,\mathbb{C})$-metabelian representations, and $SL(2,\mathbb{Z}_{3})$-representations of $\pi_{1}(S^{4}\setminus K^{m,n})$, respectively. Especially, it is proved in [14] that, for two branched twist spins $K_{1}^{m_{1},n_{1}}$ and $K_{2}^{m_{2},n_{2}}$ constructed from two $1$-knots $K_{1}$ and $K_{2}$, if $m_{1}\neq m_{2}$ then $K_{1}^{m_{1},n_{1}}$ and $K_{2}^{m_{2},n_{2}}$ are not equivalent. Note that $K^{m,n}$ and $K^{m,n+2m}$ are equivalent since $K^{m,n+m}$ is obtained from $K^{m,n}$ by the Gluck surgery along $K^{m,n}$, see the explanation in the proof of [13, Corollary 3.3]. By reversing the orientation of $S^{4}$ if necessary and applying the isomorphism between $K^{m,n}$ and $K^{m,n+2m}$, we can assume that $(m,n)$ is either a pair of copirme positive integers with $n<2m$ or $(0,1)$.
@@
To prove the above theorems, we use known results about $3$-orbifolds of cyclic type in [34, 2, 36]. The point is that the quotient group of $\pi_{1}(S^{4}\setminus K^{m,n})$ by its center can be regarded as the fundamental group of a $3$-orbifold of cyclic type unless $K$ is a torus knot (Lemma 3.1). We also use the fact that the fiber of $K^{m,n}$ is obtained as the $m$-fold cyclic branched cover of $S^{3}$ along $K$ with removing an open $3$-ball [39, 26] and its fundamental group is the commutator subgroup of $\pi_{1}(S^{4}\setminus K^{m,n})$.
@@
They would like to thank Jonathan A. Hillman, Kokoro Tanaka and Yuta Taniguchi for their helpful suggestions and precious comments. The second author is supported by JSPS KAKENHI Grant Numbers JP19K03499, JP23K03098 and JP23H00081 and JSPS-VAST Joint Research Program, Grant number JPJSBP120219602.
@@ 2.1. Branched twist spins
It will be proved in Lemma 3.1 that if $K$ is neither a trivial knot nor a torus knot then the center of $\pi_{1}(S^{4}\setminus K^{m.n})$ is $\langle h\rangle$. In this case, the quotient group of the group (2.1) by its center is
(2.2) $\pi_{1}(S^{4}\setminus K^{m,n})/\langle h\rangle\cong\langle x_{1},\ldots,x_{l}\mid r_{1},\ldots,r_{l},x_{1}^{m}\rangle.$
@@ 2.2. $3$-orbifolds
An orbifold is a connected and separable metric space locally homeomorphic to the quotient space of $\mathbb{R}^{n}$ by a finite group action. This notion is introduced by Satake [31] and renamed by W. Thurston [37, Chapter 13]. A $3$-dimensional orbifold ($3$-orbifold for short) whose ramification locus is the disjoint union of simple closed curves is called a $3$-orbifold of cyclic type in [8, 2] and called a $3$-branchfold in [33]. A $3$-orbifold is said to be sufficiently large (or Haken) if there exists an orientable and incompressible $2$-suborbifold that is not boundary parallel. See for instance [37, 34, 35, 36, 2, 3, 4] for basics of $3$-orbifolds.
@@
Note that if $\mathcal{O}(K,r)$ is sufficiently large for some $r\geq 2$ then it is for any $r\geq 2$. A sufficient condition for a prime link $L$ to satisfy that $\mathcal{O}(L,r)$ is sufficiently large for some $r\geq 2$ is given in [34, Proposition 8.2].
@@ 3. Proofs
Suppose $m\geq 2$. If $K$ is non-trivial and $K\not\in\mathcal{T}$, then the center is isomorphic to $\mathbb{Z}$, which is generated by $h$ in (2.1), and, in particular,
@@
Suppose $K$ is non-trivial. Since $h$ in (2.1) is a non-trivial element of the center of $\pi_{1}(S^{4}\setminus K^{m,n})$, $\langle h\rangle\cong\mathbb{Z}$ is a subgroup of the center of $\pi_{1}(S^{4}\setminus K^{m,n})$. An element of the center of $\pi_{1}(S^{4}\setminus K^{m,n})$ has the form $xh^{\delta}$, where $x$ is a product of the generators $x_{1},\ldots,x_{l}$ in (2.1) and $\delta\in\mathbb{Z}$. In particular, if $xh^{\delta}$ is in the center then $x$ is also. It is known that $\pi_{1}(S^{3}\setminus K)$ has a non-trivial center if and only if $K\in\mathcal{T}$ [7]. Therefore, if $K\not\in\mathcal{T}$, then $x$ should be trivial and the center of $\pi_{1}(S^{4}\setminus K^{m,n})$ is exactly $\langle h\rangle$. ∎
Lemma 3.2.
Suppose $m\geq 3$ and $K\in\mathcal{H}$ and if $m=3$ then $K$ is not a figure-eight knot. Then $\pi_{1}^{orb}(\mathcal{O}(K,m))$ does not contain a subgroup $\mathbb{Z}\oplus\mathbb{Z}$.
Proof.
By Theorem 2.3, $M_{m}(K)$ is hyperbolic. In particular, $\pi_{1}(M_{m}(K))$ does not contain a subgroup $\mathbb{Z}\oplus\mathbb{Z}$. Assume that $\pi_{1}^{orb}(\mathcal{O}(K,m))$ contains a subgroup $\mathbb{Z}\oplus\mathbb{Z}$ and let $(a,b)$ be generators of $\mathbb{Z}\oplus\mathbb{Z}$, i.e., $\mathbb{Z}\oplus\mathbb{Z}=\langle a,b\mid aba^{-1}b^{-1}\rangle$. Let $G$ be the subgroup generated by $(ma,mb)$, which is isomorphic to $\mathbb{Z}\oplus\mathbb{Z}$. The covering map $h:M_{m}(K)\to\mathcal{O}(K,m)$ induces a monomorphism $h_{*}:\pi_{1}(M_{m}(K))\to\pi_{1}^{orb}(\mathcal{O}(K,m))$. Hence, there exists a subgroup $H$ of $\pi_{1}(M_{m}(K))$ such that $h_{*}|_{H}:H\to G$ is an isomorphism. Thus $\pi_{1}(M_{m}(K))$ contains $\mathbb{Z}\oplus\mathbb{Z}$ as a subgroup, which is a contradiction. ∎
@@
If $K_{2}\in\mathcal{T}$, then we can easily confirm from (2.1) that the quotient group of $\pi_{1}(S^{4}\setminus K_{2}^{m,n})$ by its center is not trivial. Hence, by Lemma 3.1, $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are not equivalent. Suppose that $K_{2}$ is non-trivial and $K_{2}\not\in\mathcal{T}$. If $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent then $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ and $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ are isomorphic by Lemma 3.1. This implies that $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))\cong\pi_{1}^{orb}(\mathcal{O}(K_{1},m))\cong\mathbb{Z}/m\mathbb{Z}$. Let $h:M_{m}(K_{2})\to\mathcal{O}(K_{2},m)$ be the covering map. The induced map $h_{*}:\pi_{1}(M_{m}(K_{2}))\to\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ is a monomorphism. If a non-trivial element $[\gamma]$ of $\mathbb{Z}/m\mathbb{Z}$ is in the image of $h_{*}$, the preimage $h^{-1}(\gamma)$ of its representative $\gamma$ cannot be a closed curve in $M_{m}(K_{2})$. Hence, we have $h_{*}(\pi_{1}(M_{m}(K_{2})))=1$. Since $h_{*}$ is a monomorphism, $\pi_{1}(M_{m}(K_{2}))=1$. Thus, by the positive solution to the Poincaré conjecture [27, 28, 29], $M_{m}(K_{2})$ is $S^{3}$. However, since $M_{m}(K_{2})$ is the $m$-fold cyclic branched cover of $S^{3}$ along a non-trivial knot $K_{2}$, it cannot be homeomorphic to $S^{3}$ due to the positive solution to the Smith conjecture [25]. This is a contradiction. ∎
@@
Suppose that $m\geq 3$ and if $m=3$ then $K_{1}$ is not a figure-eight knot. If $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent then $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ and $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ are isomorphic by Lemma 3.1, and hence $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ also contains a subgroup $\mathbb{Z}\oplus\mathbb{Z}$. By Lemma 3.2, this is a contradiction.
@@
Due to Propositions 3.3, 3.4 and 3.5, it is enough to show the case where $K_{1},K_{2}\in\mathcal{H}$. Suppose that $K_{1},K_{2}\in\mathcal{H}$. If $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent then $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ and $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ are isomorphic by Lemma 3.1. If either $m\geq 4$ or both of $K_{1}$ and $K_{2}$ are not a figure-eight knot, then, by Theorem 2.3, $M_{m}(K_{1})$ and $M_{m}(K_{2})$ are hyperbolic and the assertion follows from the Mostow rigidity (cf. [2, Theorem 6.9]). Suppose that $m=3$, $K_{1}$ is a figure-eight knot and $K_{2}$ is not. The Hantzsche-Wendt manifold $M_{3}(K_{1})$ admits a riemannian metric with sectional curvature being constantly zero [38, 22]. Therefore, $\pi_{1}(M_{3}(K_{1}))$ is not isomorphic to $\pi_{1}(M_{3}(K_{2}))$ since $M_{3}(K_{2})$ is hyperbolic by Theorem 2.3. Hence $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are not equivalent. ∎
@@
Suppose that $K_{2}\in\mathcal{H}$. Assume that if $m=3$ then $K_{2}$ is not a figure-eight knot. For each $i=1,2$, the fiber of $K_{i}^{m,n}$ is obtained from $M_{m}(K_{i})$ by removing an open $3$-ball. Since $M_{m}(K_{1})$ is Seifert fibered and $M_{m}(K_{2})$ is hyperbolic, $\pi_{1}(M_{m}(K_{1}))$ and $\pi_{1}(M_{m}(K_{2}))$ are not isomorphic. Hence $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are not equivalent. Suppose that $m=3$ and $K_{2}$ is a figure-eight knot. It is known that the Hantzsche-Wendt manifold $M_{3}(K_{2})$ satisfies $H_{1}(M_{3}(K_{2}))=(\mathbb{Z}/4\mathbb{Z})\oplus(\mathbb{Z}/4\mathbb{Z})$, see [22, 17]. On the other hand, $M_{3}(K_{1})$ is the Brieskorn manifold of the singularity of the complex polynomial map $f(x,y,z)=x^{p}+y^{q}+z^{3}$ at the origin. Using the formulas in [23, Corollary 6.2 and Theorem 7.2], we can confirm that $|H_{1}(M_{3}(K_{1}))|$ cannot be $16$. Thus, $\pi_{1}(M_{3}(K_{2}))$ is not isomorphic to $\pi_{1}(M_{3}(K_{1}))$. Hence $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are not equivalent. ∎
@@
Suppose that $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent. Since $K_{1}$ and $K_{2}$ are prime, non-trivial and non-torus knots and $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent, $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ and $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ are isomorphic by Lemma 3.1.
For $i=1,2$, if $K_{i}$ is a prime, satellite knot then $O(K_{i},r)$ contains an incompressible torus as mentioned in the proof of Proposition 3.4, and hence $\mathcal{O}(K_{i},r)$ is sufficiently large for $r\geq 2$. If $K_{i}$ is Conway reducible, then $\mathcal{O}(K_{i},r)$ is also sufficiently large for $r\geq 2$. Hence, since $\pi_{1}(\mathcal{O}(K_{1},m))$ and $\pi_{1}(\mathcal{O}(K_{2},m))$ are isomorphic, $K_{1}$ and $K_{2}$ are equivalent by Theorem 2.2. ∎
@@
Knots without essential tori and essential Conway spheres are excluded in Theorem 1.2. It is a difficult problem to determine the class of knots determined by their double branched covers, see [21, Problem 3.25]. There are several knots that are not determined by their double branched covers, see for instance [20]. It is well-known that a mutation pair of non-equivalent Montesinos knots $K_{1}$ and $K_{2}$ has the same double branched covers. However, $\pi_{1}(\mathcal{O}(K_{1},2))$ and $\pi_{1}(\mathcal{O}(K_{2},2))$ are different by Theorem 2.2. This means that the distinction by $3$-orbifold groups works for such a pair though that by double branched covers does not.

### v4

@@ 1. Introduction
Circle actions on homotopy $4$-spheres are classified by Montgomery and Yang [28] and Fintushel [11]. Later, Pao proved in [31] that such homotopy $4$-spheres are the standard $4$-sphere $S^{4}$. A $2$-sphere embedded in $S^{4}$ that is invariant under a circle action is called a branched twist spin (cf. [20, §16.3]). Note that a circle embedded in $S^{3}$ that is invariant under a circle action is called a torus knot. A twist spun knot is an example of a branched twist spin. See Section 2.1 for more precise explanation. In this sense, branched twist spins are the simplest class of $2$-knots in $S^{4}$ if one wants to understand $2$-knots geometrically. The orbit space of a circle action on $S^{4}$ is the $3$-sphere $S^{3}$ except a trivial case and the image of the union of singular orbits by the orbit map constitutes an embedded circle, that is a $1$-knot, in $S^{3}$. Conversely, we may recover a branched twist spin from a $1$-knot in $S^{3}$ with a pair of coprime integers $m$ and $n$ with $n>0$ uniquely. For a given knot $K$ in $S^{3}$ and these integers $m$ and $n$, we denote the corresponding branched twist spin by $K^{m,n}$.
@@
The aim of this paper is to understand if we can obtain the same branched twist spin from different $1$-knots in $S^{3}$. This problem is studied by the first author in [12, 13, 15] using knot determinants, irreducible $SL(2,\mathbb{C})$-metabelian representations, and $SL(2,\mathbb{Z}_{3})$-representations of $\pi_{1}(S^{4}\setminus K^{m,n})$, respectively. Especially, it is proved in [15] that, for two branched twist spins $K_{1}^{m_{1},n_{1}}$ and $K_{2}^{m_{2},n_{2}}$ constructed from two $1$-knots $K_{1}$ and $K_{2}$, if $m_{1}\neq m_{2}$ then $K_{1}^{m_{1},n_{1}}$ and $K_{2}^{m_{2},n_{2}}$ are not equivalent. Note that $K^{m,n}$ and $K^{m,n+2m}$ are equivalent since $K^{m,n+m}$ is obtained from $K^{m,n}$ by the Gluck surgery along $K^{m,n}$, see the explanation in the proof of [14, Corollary 3.3]. By reversing the orientation of $S^{4}$ if necessary and applying the isomorphism between $K^{m,n}$ and $K^{m,n+2m}$, we can assume that $(m,n)$ is either a pair of coprime positive integers with $n<m$, $(0,1)$ or $(1,1)$.
@@
To prove the above theorems, we use known results about $3$-orbifolds of cyclic type in [39, 2]. The point is that the quotient group of $\pi_{1}(S^{4}\setminus K^{m,n})$ by its center can be regarded as the fundamental group of a $3$-orbifold of cyclic type in most cases (Lemma 3.1). We also use the fact that the fiber of $K^{m,n}$ is obtained as the $m$-fold cyclic branched cover of $S^{3}$ along $K$ with removing an open $3$-ball [44, 31] and its fundamental group is the commutator subgroup of $\pi_{1}(S^{4}\setminus K^{m,n})$.
@@
The authors would like to express their sincere gratitude to Makoto Sakuma for pointing out a gap in the proof of Lemma 3.1 and telling us how to fill it. They would also like to thank Jonathan A. Hillman, Yuya Koda, Kokoro Tanaka and Yuta Taniguchi for their helpful suggestions and precious comments. The second author is supported by JSPS KAKENHI Grant Numbers JP19K03499, JP23K03098 and JP23H00081, JSPS-VAST Joint Research Program, Grant number JPJSBP120219602, and Keio University Academic Development Funds for Individual Research.
@@ 2.1. Branched twist spins
In [31], Pao constructed the complement of $K^{m,n}$ by setting the $m$-fold cyclic branched cover of $S^{3}$ along $K$ with removing an open ball to be a fiber and the monodromy to be periodic, shifting the sheets of the cyclic branched cover by $n$, and then proved that $K^{m,n}$ is a $2$-knot in $S^{4}$. His results are summarized as follows:
Theorem 2.2 (Pao [31]).
If $m\geq 1$, then $K^{m,n}$ is a fibered 2-knot in $S^{4}$ whose fiber is diffeomorphic to the $m$-fold cyclic branched cover of $S^{3}$ along $K$ with removing an open ball and whose monodromy is periodic, shifting the sheets of the cyclic branched cover by $n$.
@@
It will be proved in Lemma 3.1 that if $K$ is neither a trivial knot nor a torus knot and the $m$-fold cyclic branched cover of $S^{3}$ along $K$ is aspherical, then the center of $\pi_{1}(S^{4}\setminus K^{m.n})$ is $\langle h\rangle$. In this case, the quotient group of the group (2.1) by its center is
$\pi_{1}(S^{4}\setminus K^{m,n})/\langle h\rangle\cong\langle x_{1},\ldots,x_{l}\mid r_{1},\ldots,r_{l},x_{1}^{m}\rangle.$
@@ 2.2. $3$-orbifolds
An orbifold is a connected and separable metric space locally homeomorphic to the quotient space of $\mathbb{R}^{n}$ by a finite group action. This notion is introduced by Satake [36] and renamed by W. Thurston [42, Chapter 13]. A $3$-dimensional orbifold ($3$-orbifold for short) whose ramification locus is the disjoint union of simple closed curves is called a $3$-orbifold of cyclic type in [9, 2] and called a $3$-branchfold in [38]. A $3$-orbifold is said to be sufficiently large (or Haken) if there exists an orientable and incompressible $2$-suborbifold that is not boundary parallel. See for instance [42, 39, 40, 41, 2, 3, 4] for basics of $3$-orbifolds.
@@
Note that if $\mathcal{O}(K,r)$ is sufficiently large for some $r\geq 2$ then it is for any $r\geq 2$. A sufficient condition for a prime link $L$ to satisfy that $\mathcal{O}(L,r)$ is sufficiently large for some $r\geq 2$ is given in [39, Proposition 8.2].
@@ 3. Proofs
Suppose $m\geq 2$. If $K$ is non-trivial and $K\not\in\mathcal{T}$ and $M_{m}(K)$ is aspherical, then the center of $\pi_{1}(S^{4}\setminus K^{m,n})$ is isomorphic to $\mathbb{Z}$, which is generated by $h$ in (2.1), and, in particular,
@@
Let $K$ be a non-trivial knot and assume that $K\not\in\mathcal{T}$ and $M_{m}(K)$ is aspherical. First we prove that the center of $\pi_{1}^{orb}(\mathcal{O}(K,m))$ is trivial.
[…]
Assume that there exists a non-trivial element $\gamma$ in the center of $\pi_{1}^{orb}(\mathcal{O}(K,m))$ of infinite order. Then $\gamma^{m}$ is a non-trivial element in the center of $\pi_{1}(M_{m}(K))$. Hence, by the positive answer to the Seifert fiber space conjecture [16, 7] (cf. [30]), $M_{m}(K)$ is a Seifert fibered space. By [26], the action of $\mathbb{Z}/m\mathbb{Z}$ on $M_{m}(K)$ is conjugate to a fiber-preserving action. Since the action of $\mathbb{Z}/m\mathbb{Z}$ on $\pi_{1}(M_{m}(K))$ sends $\gamma^{m}$ to itself, the action of $\mathbb{Z}/m\mathbb{Z}$ on $M_{m}(K)$ preserves the orientations of the fibers. This implies that a fiber in $M_{m}(K)$ intersecting the preimage $\tilde{K}$ of $K$ should coincide with $\tilde{K}$ itself since $\tilde{K}$ is the fixed point set of the action of $\mathbb{Z}/m\mathbb{Z}$. Thus, $\mathcal{O}(K,m)=M_{m}(K)/(\mathbb{Z}/m\mathbb{Z})$ is a Seifert fibered orbifold consisting of only circle fibers one of which is $K$. In particular, $\mathcal{O}(K,m)\setminus K$ is a Seifert fibered space. This contradicts the assumption that $K\not\in\mathcal{T}$.
Thus, we conclude that the center of $\pi_{1}^{orb}(\mathcal{O}(K,m))$ is trivial.
Set $G=\pi_{1}(S^{4}\setminus K^{m,n})$ and let $Z(G)$ denote the center of $G$. By (2.1), the subgroup of $G$ generated by $h$ is a subgroup of $Z(G)$. Each element in $Z(G)$ is written in the form $xh^{\delta}$, where $x$ is a product of the generators $x_{1},\ldots,x_{l}$ in (2.1) and $\delta\in\mathbb{Z}$. In particular, if $xh^{\delta}$ is in $Z(G)$ then $x$ is also.
Let $y$ be a non-trivial element in $Z(G)$ that is given as a product of the generators $x_{1},\ldots,x_{l}$ in (2.1) but cannot be presented as a product of $x_{1}^{m},\ldots,x_{l}^{m}$. Assume that $yx_{i}y^{-1}x_{i}^{-1}$ is trivial in $G$ for $i=1,\ldots,l$. Then, there exists a sequence of Tietze transformations that changes $yx_{i}y^{-1}x_{i}^{-1}$ to $1$, possibly including the word $h$ in the middle. Since $h\in Z(G)$, we obtain a sequence of Tietze transformations that changes $yx_{i}y^{-1}x_{i}^{-1}$ to $1$, without $h$, up to adding and deleting the words $x_{1}^{m},\ldots,x_{l}^{m}$ anywhere in the middle. This means that $yx_{i}y^{-1}x_{i}^{-1}$ is trivial in the orbifold group $\pi_{1}^{orb}(\mathcal{O}(K,m))$ for all $i=1,\ldots,l$, that is, $y$ is in the center of $\pi_{1}^{orb}(\mathcal{O}(K,m))$. Since the center of $\pi_{1}^{orb}(\mathcal{O}(K,m))$ is trivial, $y$ is trivial in $\pi_{1}^{orb}(\mathcal{O}(K,m))$. This contradicts the assumption that $y$ is non-trivial.
Now we see that any element in $Z(G)$ given as a product of the generators $x_{1},\ldots,x_{l}$ is presented as a product of $x_{1}^{m},\ldots,x_{l}^{m}$. This means that $Z(G)$ is generated by $h$. Setting $h$ in the presentation (2.1) to be trivial, we obtain $G/Z(G)\cong\langle x_{1},\ldots,x_{l}\mid r_{1},\ldots,r_{l},\,x_{1}^{m}\rangle$. Thus the assertion holds. ∎
@@
Assume that $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent. Then $\pi_{1}(S^{4}\setminus K_{1}^{m,n})$ and $\pi_{1}(S^{4}\setminus K_{2}^{m,n})$ are isomorphic and hence their commutator subgroups are also isomorphic. The fiber of $K_{i}^{m,n}$, for $i=1,2$, is the $m$-fold cyclic branched cover $M_{m}(K_{i})$ of $S^{3}$ along $K_{i}$ with removing an open $3$-ball, and hence the commutator subgroup of $\pi_{1}(S^{4}\setminus K_{i}^{m,n})$ is isomorphic to $\pi_{1}(M_{m}(K_{i}))$. Thus $\pi_{1}(M_{m}(K_{1}))$ and $\pi_{1}(M_{m}(K_{2}))$ are isomorphic. Since $M_{m}(K_{1})$ is $S^{3}$, $M_{m}(K_{2})$ is also by the positive solution to the Poincaré conjecture [32, 33, 34]. However, since $M_{m}(K_{2})$ is the $m$-fold cyclic branched cover of $S^{3}$ along a non-trivial knot $K_{2}$, it cannot be homeomorphic to $S^{3}$ due to the positive solution to the Smith conjecture [29]. This is a contradiction. ∎
@@
Suppose that $m\geq 3$ and if $m=3$ then $K_{1}$ is not a figure-eight knot. If $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent then $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ and $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ are isomorphic by Lemma 3.1, and hence $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ also contains a subgroup $\mathbb{Z}\oplus\mathbb{Z}$. However, the orbifold $\mathcal{O}(K_{1},m)$ is hyperbolic and compact by Theorem 2.4. Hence $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ does not contain a subgroup $\mathbb{Z}\oplus\mathbb{Z}$. This is a contradiction.
@@
Due to Propositions 3.2, 3.3 and 3.4, it is enough to show the case where $K_{1},K_{2}\in\mathcal{H}$. Suppose that $K_{1},K_{2}\in\mathcal{H}$. If $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent then $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ and $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ are isomorphic by Lemma 3.1. If either $m\geq 4$ or both of $K_{1}$ and $K_{2}$ are not a figure-eight knot, then, by Theorem 2.4, $M_{m}(K_{1})$ and $M_{m}(K_{2})$ are hyperbolic and the assertion follows from the Mostow rigidity (cf. [3, Theorem 6.9]). Suppose that $m=3$, $K_{1}$ is a figure-eight knot and $K_{2}$ is not. The Hantzsche-Wendt manifold $M_{3}(K_{1})$ admits a riemannian metric with sectional curvature being constantly zero [43, 24]. Therefore, $\pi_{1}(M_{3}(K_{1}))$ is not isomorphic to $\pi_{1}(M_{3}(K_{2}))$ since $M_{3}(K_{2})$ is hyperbolic by Theorem 2.4. Hence $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are not equivalent. ∎
@@
Suppose that $K_{2}\in\mathcal{H}$. Assume that if $m=3$ then $K_{2}$ is not a figure-eight knot. For each $i=1,2$, the fiber of $K_{i}^{m,n}$ is obtained from $M_{m}(K_{i})$ by removing an open $3$-ball. Since $M_{m}(K_{1})$ is Seifert fibered and $M_{m}(K_{2})$ is hyperbolic, $\pi_{1}(M_{m}(K_{1}))$ and $\pi_{1}(M_{m}(K_{2}))$ are not isomorphic. Hence $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are not equivalent. Suppose that $m=3$ and $K_{2}$ is a figure-eight knot. It is known that the Hantzsche-Wendt manifold $M_{3}(K_{2})$ satisfies $H_{1}(M_{3}(K_{2}))=(\mathbb{Z}/4\mathbb{Z})\oplus(\mathbb{Z}/4\mathbb{Z})$, see [24, 19]. On the other hand, $M_{3}(K_{1})$ is the Brieskorn manifold of the singularity of the complex polynomial map $f(x,y,z)=x^{p}+y^{q}+z^{3}$ at the origin. Using the formulas in [25, Corollary 6.2 and Theorem 7.2], we can confirm that $|H_{1}(M_{3}(K_{1}))|$ cannot be $16$. Thus, $\pi_{1}(M_{3}(K_{2}))$ is not isomorphic to $\pi_{1}(M_{3}(K_{1}))$. Hence $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are not equivalent. ∎
@@
Suppose that $K_{1}^{m,n}$ and $K_{2}^{m,n}$ are equivalent. Since $K_{1}$ and $K_{2}$ are prime, non-trivial and non-torus knots and $M_{m}(K_{1})$ and $M_{m}(K_{2})$ are aspherical, $\pi_{1}^{orb}(\mathcal{O}(K_{1},m))$ and $\pi_{1}^{orb}(\mathcal{O}(K_{2},m))$ are isomorphic by Lemma 3.1.
For $i=1,2$, if $K_{i}$ is a prime, satellite knot then $\mathcal{O}(K_{i},r)$ contains an incompressible torus as mentioned in the proof of Proposition 3.3, and hence $\mathcal{O}(K_{i},r)$ is sufficiently large for $r\geq 2$. If $K_{i}$ is Conway reducible, then $\mathcal{O}(K_{i},r)$ is also sufficiently large for $r\geq 2$. Hence, since $\pi_{1}(\mathcal{O}(K_{1},m))$ and $\pi_{1}(\mathcal{O}(K_{2},m))$ are isomorphic, $K_{1}$ and $K_{2}$ are equivalent by Theorem 2.3. ∎
@@
Knots without essential tori and essential Conway spheres are excluded in Theorem 1.2. It is a difficult problem to determine the class of knots determined by their double branched covers, see [23, Problem 3.25]. There are several knots that are not determined by their double branched covers, see for instance [22]. It is well-known that a mutation pair of non-equivalent Montesinos knots $K_{1}$ and $K_{2}$ has the same double branched covers. However, $\pi_{1}(\mathcal{O}(K_{1},2))$ and $\pi_{1}(\mathcal{O}(K_{2},2))$ are different by Theorem 2.3. This means that the distinction by $3$-orbifold groups works for such a pair though that by double branched covers does not.